OpenStudy (smurfy14):

ln2000=ln20+0.825t what is the answer and how do you get it?

myininaya (myininaya):

put the natural logs on one side and then combine them using lnx/y=lnx-lny property and then base e both sides

OpenStudy (anonymous):

ln 2000 - ln 20 = 0.825t ln (2000/20) = 0.825 t t = (ln(2000/20)) / 0.825

myininaya (myininaya):

i mean no base e

myininaya (myininaya):

lol

OpenStudy (smurfy14):

so it goes like this: ln200/ln20=.825t ln100=.825t t=5.582?

myininaya (myininaya):

no it goes like this first step subtract ln20 on both sides giving you: ln2000-ln20=.0825t

myininaya (myininaya):

now we can write ln[(2000/20)]=.0825t

myininaya (myininaya):

ln(100)=.0825t

myininaya (myininaya):

to solve for t we just divide now by .0825giving us [ln(100)]/(.0825)=t

myininaya (myininaya):

remember that property: ln(x/y)=lnx-lny?

OpenStudy (smurfy14):

but isnt that what i did? im confused...you said to ln2000-ln20 (subtract) but instead you divided?

OpenStudy (smurfy14):

so what would be the final answer?

myininaya (myininaya):

thats how i would write my final answer. ln(x/y) does not equal lnx/lny

OpenStudy (smurfy14):

oh ok i see so final answer when put in calc would be??

myininaya (myininaya):

> ln(100.0)/(.0825); 55.82024467 >